OpenStudy (anonymous):

Help i'm a 6th grader needing help

OpenStudy (anonymous):

this is the file

OpenStudy (anonymous):

this is the file

OpenStudy (anonymous):

What file?

OpenStudy (anonymous):

never mind for some reason it forbiden

OpenStudy (anonymous):

try refreshing your page

OpenStudy (anonymous):

Did it work?

OpenStudy (anonymous):

ill just ask u the question :3

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

find the factorization of 36. write the prime factorization using exponents

OpenStudy (anonymous):

ok so what is all the factors of 36?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

1, 2, 3, 4, 6, 9, 12, 18, 36

OpenStudy (anonymous):

oh hi

OpenStudy (misty1212):

prime factorization right?

OpenStudy (anonymous):

ya i have to right it with exponents

OpenStudy (misty1212):

some people do this with trees, i do not \[36=2\times 18\] is a start then since \(18=2\times 9\) you have \[36=2^2\times 9\]

OpenStudy (misty1212):

then since \(9=3^2\)the prime factorization is \[36=2^2\times 3^2\]

OpenStudy (anonymous):

thx you have been really helpful can u teach me how to do this method

OpenStudy (misty1212):

keep dividing by the smallest number you can think of that divides evenly them move on to the next one

OpenStudy (anonymous):

thx i know dis is going to help me in the future thx a whole bunch

OpenStudy (misty1212):

\[\color\magenta\heartsuit \]